OpenStudy (anonymous):

factor completely 4x^2 - 24x - 8 This is just simple factoring. I believe it may be prime, because I can only get as far as factoring out the 4. I end up with 4(x^2 - 6x - 8). I am unable to factor what's inside the parenthesis, so I am assuming the expression is Prime.

OpenStudy (anonymous):

You can use the X method. Are you familiar with that?

OpenStudy (imstuck):

Use the quadratic formula. It's never-fail. Do you know it?

OpenStudy (anonymous):

Heres a cool song if you want to remember it. https://www.youtube.com/watch?v=H_7lNT9oDzI

OpenStudy (anonymous):

no I don't know it, and what does a song have anything to do with anything?

OpenStudy (anonymous):

oh im sorry i replied to the wrong thread. someone else is working with the quadratic formula

OpenStudy (imstuck):

The song helps your remember the quadratic formula. It's quite something, all the variables you have to remember.\[\frac{ -b \pm \sqrt{(b)^{2}-4ac} }{2a }\]If you can remember all that then maybe you don't need the song.

OpenStudy (imstuck):

Anyways, factor your expression inside the parenthesis using the quadratic formula and the fact that in your expression, a = 1, b = -6, c = -8\[\frac{ -(-6)\pm \sqrt{(-6)^{2}-4(1)(-8)} }{ 2(1) }\]

OpenStudy (imstuck):

That gives you\[\frac{ 6\pm \sqrt{36+32} }{2 }\]or\[\frac{ 6+\sqrt{68} }{ 2 } , \frac{ 6-\sqrt{68} }{ 2 }\]

OpenStudy (anonymous):

Yep thats right ^. lol IMStuck how much time do you spend on here. you helped me once too.

OpenStudy (imstuck):

The first one gives you x = 7.123 and the other one gives you x = -1.123

OpenStudy (imstuck):

I spend literally almost all day on here. I'm a teacher (going through withdrawals) and want to keep up with my math all year long. I tutor on tutor.com, too! I love love math!

OpenStudy (anonymous):

I am not trying to solve an equation. I just need to factor the problem. The answer should be in the form of ( )( ). This is why I think it is prime.

OpenStudy (imstuck):

It can factor into (x-7.123)(x+1.123). It's not prime.

OpenStudy (imstuck):

No polynomial is EVER unfactorable, ever!!!! The quadratic formula can factor any second degree polynomial ALWAYS.

OpenStudy (imstuck):

...trust me...

OpenStudy (anonymous):

You can factor out a 4.

OpenStudy (anonymous):

can not be factored

OpenStudy (imstuck):

The quadratic formula is used for factoring.

OpenStudy (anonymous):

\(4x^2 - 24x - 8\) \(= 4(x^2 - 6x - 2)\) Now it's factored completely.

OpenStudy (anonymous):

Thank you. We are not that far in the class to use the quadratic formula and there have problems where the answer has been prime, so I just thought this one was prime. At this point in the class I believe it has to equal the exact number in the polynomial. For example, 7.123 * 1.123 = 7.999. Thanks MaxwellSmart...I wasn't sure if that was considered to me completely factored or not.

OpenStudy (anonymous):

You're welcome.